# A Survey on Large Language Models for Code Generation

### v1

@@ 4.6. Prompting Engineering
Large-scale language models (LLMs) such as GPT-3 and its successors have been trained on large-scale data corpora, endowing them with substantial world knowledge (Brown et al., 2020; Wei et al., 2021; Ouyang et al., 2022). Despite this, crafting an effective prompt to harness the full potential of LLMs remains a long-standing challenge (Liu et al., 2023c). Recent advancements in prompting engineering have expanded the capabilities of LLMs, enabling more sophisticated task completion and enhancing both reliability and performance. Notable techniques include Chain-of-Thought (CoT) (Wei et al., 2022b), Self-Consistency (Wang et al., 2022b), Tree-of-Thought (ToT) (Yao et al., 2024), Reasoning via Planning (RAP) (Hao et al., 2023), ReAct (Yao et al., 2023), Self-Refine (Madaan et al., 2024), Reflexion (Shinn et al., 2024), and LATS (Zhou et al., 2023).
@@
Prompting engineering is particularly advantageous as it bypasses the need for additional training and can significantly elevate performance. Consequently, numerous studies have leveraged this technique for iterative and self-improving (refining) code generation within proprietary LLMs such as ChatGPT and GPT-4. Figure 7 illustrates the general pipeline for self-improving code generation with LLMs. For instance, Self-Debugging (Chen et al., 2023) involves prompting an LLM to iteratively refine a predicted program by utilizing feedback composed of code explanations combined with execution results, which assists in identifying and rectifying errors. When unit tests are unavailable, this feedback can rely solely on code explanations. In parallel, SelfEvolve (Jiang et al., 2023) employs a two-stage process where LLMs first generate domain-specific knowledge for a problem, followed by a trial code. This code is then iteratively refined through interactive prompting and execution feedback. An empirical investigation by (Olausson et al., 2023) provides a comprehensive analysis of the self-repairing capabilities for code generation in models like Code Llama, GPT-3.5, and GPT-4, using problem sets from HumanEval and APPS. This study yields a series of insightful observations and findings, shedding light on the self-refinement effectiveness of these LLMs. Moreover, Reflexion (Shinn et al., 2024) introduces a general approach for code generation wherein LLM-powered agents engage in verbal self-reflection on task feedback signals, storing these reflections in an episodic memory buffer to inform and improve decision-making in subsequent interactions. LATS (Zhou et al., 2023) adopts a novel strategy, utilizing LLMs as agents, value functions, and optimizers. It enhances decision-making by meticulously constructing trajectories through Monte Carlo Tree Search (MCTS) algorithms, integrating external feedback, and learning from experience. This approach has demonstrated remarkable results in code generation, achieving a pass@1 of 94.4% on the HumanEval benchmark with GPT-4.
@@
Figure 7. An illustration of the self-improving code generation pipeline using prompts for large language models (LLMs). This process incorporates iterative self-refinement by integrating execution outcomes and includes an optional self-reflection mechanism to enhance generation quality.
@@
Distinct from the aforementioned methods, CodeT (Chen et al., 2022b) and LEVER (Ni et al., 2023a) prompt LLMs to generate numerous code samples, which are then re-ranked based on execution outcomes to select the optimal solution. Notably, these approaches do not incorporate a self-refinement step to further improve code generation.
@@ 4.7. Repository Level & Long Context
In contemporary software engineering practices, modifications to a code repository are widespread and encompass a range of activities, including package migration, temporary code edits, and the resolution of GitHub issues. While large language models (LLMs) showcase impressive prowess in function-level code generation, they often falter when grappling with the broader context inherent to a repository, such as import dependencies, parent classes, and files bearing similar names. These deficiencies result in suboptimal performance in repository-level code generation, as identified in recent studies (Shrivastava et al., 2023b, a). The challenges faced by LLMs in this domain are primarily due to the following factors:
@@
Accordingly, a surge of benchmarks tailored for repository-level code generation has emerged, such as RepoEval (Zhang et al., 2023c), Stack-Repo (Shrivastava et al., 2023a), Repobench (Liu et al., 2023b), EvoCodeBench (Li et al., 2024), SWE-bench (Jimenez et al., 2023), CrossCodeEval (Ding et al., 2024), and SketchEval (Zan et al., 2024). The detailed statistics and comparisons of these benchmarks are presented in Table 3.
Despite the progress made by these methods in repository-level code generation, significant challenges remain to be addressed. Programming developers are often required to invest considerable time in editing and debugging (Vaithilingam et al., 2022; Mozannar et al., 2022; Shrivastava et al., 2023a; Barke et al., 2023; Bird et al., 2022). However, the advent of LLM-powered coding agents, such as AutoCodeRover (Zhang et al., 2024), SWE-Agent (John Yang, 2024), and OpenDevin (OpenDevin, 2024), has demonstrated their potential to tackle complex problems, paving the way for future exploration in this field (for more details, see Section 4.9).
@@ 4.8. Retrieval Augmented
Large Language Models (LLMs) have exhibited impressive capabilities but are hindered by several critical issues such as hallucination (Liang et al., 2023; Zhang et al., 2023e), obsolescence of knowledge (Jang et al., 2022), and non-transparent (Bommasani et al., 2021), untraceable reasoning processes (Zhou et al., 2022b; Wei et al., 2022b; Huang and Chang, 2023; Gao et al., 2023b). While techniques like instruction-tuning (see Section 4.4) and reinforcement learning with feedback (see Section 4.5) mitigate these issues, they also introduce new challenges, such as catastrophic forgetting and the requirement for substantial computational resources during training (Ovadia et al., 2023; Gupta et al., 2024).
@@ 4.9. Autonomous Coding Agents
The advent of large language models (LLMs) has marked the beginning of a new era of potential pathways toward artificial general intelligence (AGI), capturing significant attention in both academia and industry (Xi et al., 2023; Weng, 2023; Wang et al., 2024b; Huang et al., 2024). A rapidly expanding array of applications for LLM-based autonomous agents, including AutoGPT (aut, 2023), AgentGPT (age, 2023), BabyAGI (bab, 2023), and AutoGen (Wu et al., 2023), underlines the promise of this technology.
LLM-powered autonomous agents are systems endowed with sophisticated reasoning abilities, leveraging an LLM as a central computational engine or controller. This allows them to formulate and execute problem-solving plans through a series of tool-enabled functions or API calls. Moreover, these agents are designed to function within a shared environment where they can communicate and engage in cooperative, competitive, or negotiating interactions (Huang et al., 2023; Wu et al., 2023; Wang et al., 2024b). The typical architecture of such an agent encompasses an LLM-based Agent, a memory module, a planning component, and a tool utilization module, as depicted in Figure 9.
@@
In the realm of automated code generation, LLM-powered autonomous agents have demonstrated remarkable proficiency. For instance, AgentCoder (Huang et al., 2023) achieved a groundbreaking pass@1 of 96.3% on the HumanEval benchmark, forwarding a step closer to the future of automated software development (Ishibashi and Nishimura, 2024). The innovative meta-programming framework termed MetaGPT (Hong et al., 2023) integrates human workflow efficiencies into LLM-based multi-agent collaboration. Furthermore, (Huang et al., 2023) introduces AgentCoder, a multi-agent framework composed of three specialized agents, each with distinct roles and capabilities. These roles include a programmer agent responsible for code generation, a test designer agent tasked with generating unit test cases, and a test executor agent that executes the code and provides feedback. This division of labor within AgentCoder promotes more efficient and effective code generation. CodeAct (Wang et al., 2024a) distinguishes itself by utilizing executable Python code to consolidate LLM agent actions within a unified action space, in contrast to the generation of JSON or textual formats. Additionally, AutoCodeRover (Zhang et al., 2024) is proposed to autonomously resolve GitHub issues for program enhancement.
@@
Despite these remarkable advancements, the journey toward fully realized AI software engineers employing LLM-powered autonomous agents is far from complete (Xi et al., 2023; Wang et al., 2024b). Critical aspects such as prompt design, context length, agent count, and toolsets call for further refinement and optimization, especially as problem complexities escalate (Ishibashi and Nishimura, 2024).
@@ 4.10. Evaluation
Table 6. The performance comparison of LLMs for code generation on the HumanEval (Chen et al., 2021) benchmark, measured by Pass@{1, 10, 100}. For models with various sizes, we report only the largest size version of each model.
Model Size pass@k
$k=1$ $k=10$ $k=100$
GPT-4 (Achiam et al., 2023) - 84.1 - -
GPT-3.5-Turbo (OpenAI, 2022) - 76.2 - -
Claude-3-Opus (Anthropic, 2024) - 82.9 - -
Claude-3-Haiku (Anthropic, 2024) - 76.8 - -
Claude-3-Sonnet (Anthropic, 2024) - 70.7 - -
StarCoder2-Instruct (Yuxiang Wei, 2024) 15.5B 72.6 - -
Llama3 (Meta, 2024) 70B 81.7 - -
CodeGemma (CodeGemma Team et al., 2024) 7B 44.5 - -
StarCoder 2 (Lozhkov et al., 2024) 15B 46.3 - -
phi-2 (Mojan Javaheripi, 2023) 2.7B 49.4 - -
WaveCoder (Yu et al., 2023) 6.7B 75 - -
StableCode (Pinnaparaju et al., 2024) 3B 29.3 - -
CodeShell (Xie et al., 2024) 7B 34.32 - -
CodeQwen (Team, 2024) 14B 45.1 - -
DeepSeek-Coder (Guo et al., 2024) 33B 56.1 - -
replit-code (Replit, 2023) 3B 20.12 - -
Phi-1.5 (Li et al., 2023b) 1.3B 41.4 - -
PanGu-Coder2 (Shen et al., 2023) 15B 61.64 79.55 91.75
WizardCoder (Luo et al., 2023) 15B 57.3 73.2 90.46
CodeFuse (Liu et al., 2023a) 34B 74.4 - -
Phi-1 (Gunasekar et al., 2023) 1.3B 50.6 - -
Code Llama (Roziere et al., 2023) 34B 48.8 76.8 93.0
OctoCoder (Muennighoff et al., 2023) 15.5B 46.2 - -
PaLM-Coder (Chowdhery et al., 2023) 540B 36 - 88.4
CodeGeeX2 (Zheng et al., 2023c) 6B 35.9 62.6 88.3
InstructCodeT5+ (Wang et al., 2023b) 16B 35.0 54.5 77.9
CodeGen-NL (Nijkamp et al., 2022) 16.1B 14.24 23.46 38.33
CodeGen-Multi (Nijkamp et al., 2022) 16.1B 18.32 32.07 50.8
CodeGen-Mono (Nijkamp et al., 2022) 16.1B 29.28 49.86 75
StarCoder (Li et al., 2023a) 15B 33.60 45.78 79.82
CodeT5+ (Wang et al., 2021) 16B 30.9 51.6 76.7
LLaMA2 (Touvron et al., 2023b) 70B 30.5 59.4 87.0
Codex (Chen et al., 2021) 12B 28.81 46.81 72.31
PaLM (Chowdhery et al., 2023) 540B 26.2 - 76.2
PanGu-Coder (Christopoulou et al., 2022) 2.6B 23.78 35.36 51.24
LLaMA (Touvron et al., 2023a) 65B 23.7 - 79.3
CodeGeeX (Zheng et al., 2023c) 13B 22.89 39.57 60.92
Replit (Replit, 2016) 3B 21.9 - -
CodeGen2 (Nijkamp et al., 2023) 16B 20.46 36.5 56.71
SantaCoder (Allal et al., 2023) 1.1B 18 29 49
AlphaCode (Li et al., 2022a) 1.1B 17.1 28.2 45.3
BLOOM (Le Scao et al., 2023) 176B 15.52 32.20 55.45
GPT-NeoX (Black et al., 2022) 20B 15.4 25.6 41.2
InCoder (Fried et al., 2022) 6.7B 15.2 27.8 47.0
GPT-J (Wang and Komatsuzaki, 2021) 6B 11.62 15.74 27.74
PyCodeGPT (Zan et al., 2022) 110M 8.33 13.36 19.13
GPT-Neo (Black et al., 2021) 2.7B 6.41 11.27 21.37
PolyCoder (Xu et al., 2022) 2.7B 5.59 9.84 17.68
JuPyT5 (Chandel et al., 2022) 300M 5.4 15.46 25.60
CodeParrot (Tunstall et al., 2022) 1.5B 3.99 8.69 17.88
@@
Despite the impressive capabilities of large language models (LLMs), they exhibit a range of behaviors that are both beneficial and potentially risky. These behaviors can enhance performance across various downstream tasks but may also introduce reliability and trustworthiness concerns in LLM deployment (Chen et al., 2021; Xu et al., 2022; Chang et al., 2024). Consequently, it is imperative to develop precise evaluation approaches to discern the qualitative and quantitive differences between models, thereby encouraging further advancements in LLM capabilities.

### v2

@@ 5.6. Prompting Engineering
Large-scale language models (LLMs) such as GPT-3 and its successors have been trained on large-scale data corpora, endowing them with substantial world knowledge (Brown, 2020; Wei et al., 2021; Ouyang et al., 2022). Despite this, crafting an effective prompting as a means of communicating with LLMs to harness their full potential remains a long-standing challenge (Liu et al., 2023d). Recent advancements in prompting engineering have expanded the capabilities of LLMs, enabling more sophisticated task completion and enhancing both reliability and performance. Notable techniques include Chain-of-Thought (CoT) (Wei et al., 2022b), Self-Consistency (Wang et al., 2022c), Tree-of-Thought (ToT) (Yao et al., 2024), Program of Thoughts (PoT) (Chen et al., 2022b), Reasoning via Planning (RAP) (Hao et al., 2023), ReAct (Yao et al., 2023), Self-Refine (Madaan et al., 2024), Reflexion (Shinn et al., 2024), and LATS (Zhou et al., 2023b). For instance, CoT significantly improves the LLMs’ ability to perform complex reasoning by providing a few chain-of-thought demonstrations as exemplars in prompting.
@@
Prompting engineering is particularly advantageous as it bypasses the need for additional training and can significantly elevate performance. Consequently, numerous studies have leveraged this technique for iterative and self-improving (refining) code generation within proprietary LLMs such as ChatGPT and GPT-4. Figure 12 illustrates the general pipeline for self-improving code generation with LLMs. For instance, Self-Debugging (Chen et al., 2023) involves prompting an LLM to iteratively refine a predicted program by utilizing feedback composed of code explanations combined with execution results, which assists in identifying and rectifying errors. When unit tests are unavailable, this feedback can rely solely on code explanations. Similarly, LDB (Zhong et al., 2024) prompts LLMs to refine generated code by incorporating debugging feedback, which consists of the evaluation of the correctness of variable values throughout runtime execution, as assessed by the LLMs. In parallel, SelfEvolve (Jiang et al., 2023) employs a two-stage process where LLMs first generate domain-specific knowledge for a problem, followed by a trial code. This code is then iteratively refined through interactive prompting and execution feedback. An empirical investigation by (Olausson et al., 2023) provides a comprehensive analysis of the self-repairing capabilities for code generation in models like Code Llama, GPT-3.5, and GPT-4, using problem sets from HumanEval and APPS. This study yields a series of insightful observations and findings, shedding light on the self-refinement effectiveness of these LLMs. Moreover, Reflexion (Shinn et al., 2024) introduces a general approach for code generation wherein LLM-powered agents engage in verbal self-reflection on task feedback signals, storing these reflections in an episodic memory buffer to inform and improve decision-making in subsequent interactions. LATS (Zhou et al., 2023b) adopts a novel strategy, utilizing LLMs as agents, value functions, and optimizers. It enhances decision-making by meticulously constructing trajectories through Monte Carlo Tree Search (MCTS) algorithms, integrating external feedback, and learning from experience. This approach has demonstrated remarkable results in code generation, achieving a pass@1 of 94.4% on the HumanEval benchmark with GPT-4.
@@
Figure 12. An illustration of the self-improving code generation pipeline using prompts for LLMs. This process incorporates iterative self-refinement by integrating execution outcomes and includes an optional self-reflection mechanism to enhance generation quality.
@@
Distinct from the aforementioned methods, CodeT (Chen et al., 2022c) and LEVER (Ni et al., 2023a) prompt LLMs to generate numerous code samples, which are then re-ranked based on execution outcomes to select the optimal solution. Notably, these approaches do not incorporate a self-refinement step to further improve code generation.
@@ 5.7. Repository Level & Long Context
In contemporary software engineering practices, modifications to a code repository are widespread and encompass a range of activities, including package migration, temporary code edits, and the resolution of GitHub issues. While LLMs showcase impressive prowess in function-level code generation, they often falter when grappling with the broader context inherent to a repository, such as import dependencies, parent classes, and files bearing similar names. These deficiencies result in suboptimal performance in repository-level code generation, as identified in recent studies (Shrivastava et al., 2023b, a). The challenges faced by LLMs in this domain are primarily due to the following factors:
@@
Accordingly, a surge of benchmarks tailored for repository-level code generation has emerged, such as RepoEval (Zhang et al., 2023c), Stack-Repo (Shrivastava et al., 2023a), Repobench (Liu et al., 2023c), EvoCodeBench (Li et al., 2024), SWE-bench (Jimenez et al., 2023), CrossCodeEval (Ding et al., 2024), and SketchEval (Zan et al., 2024). The detailed statistics and comparisons of these benchmarks are presented in Table 6.
Despite the progress made by these methods in repository-level code generation, significant challenges remain to be addressed. Programming developers are often required to invest considerable time in editing and debugging (Vaithilingam et al., 2022; Mozannar et al., 2022; Shrivastava et al., 2023a; Barke et al., 2023; Bird et al., 2022). However, the advent of LLM-powered coding agents, such as AutoCodeRover (Zhang et al., 2024b), SWE-Agent (John Yang, 2024), and OpenDevin (OpenDevin, 2024), has demonstrated their potential to tackle complex problems, paving the way for future exploration in this field (for more details, see Section 5.9).
@@ 5.8. Retrieval Augmented
LLMs have exhibited impressive capabilities but are hindered by several critical issues such as hallucination (Liang et al., 2023; Zhang et al., 2023e), obsolescence of knowledge (Jang et al., 2022), and non-transparent (Bommasani et al., 2021), untraceable reasoning processes (Zhou et al., 2022b; Wei et al., 2022b; Huang and Chang, 2023; Gao et al., 2023b). While techniques like instruction-tuning (see Section 5.4) and reinforcement learning with feedback (see Section 5.5) mitigate these issues, they also introduce new challenges, such as catastrophic forgetting and the requirement for substantial computational resources during training (Ovadia et al., 2023; Gupta et al., 2024).
@@ 5.9. Autonomous Coding Agents
The advent of LLMs has marked the beginning of a new era of potential pathways toward artificial general intelligence (AGI), capturing significant attention in both academia and industry (Xi et al., 2023; Weng, 2023; Wang et al., 2024c; Huang et al., 2024). A rapidly expanding array of applications for LLM-based autonomous agents, including AutoGPT (aut, 2023), AgentGPT (age, 2023), BabyAGI (bab, 2023), and AutoGen (Wu et al., 2023), underlines the promise of this technology.
LLM-powered autonomous agents are systems endowed with sophisticated reasoning abilities, leveraging an LLM as a central computational engine or controller. This allows them to formulate and execute problem-solving plans through a series of tool-enabled functions or API calls. Moreover, these agents are designed to function within a shared environment where they can communicate and engage in cooperative, competitive, or negotiating interactions (Huang et al., 2023; Wu et al., 2023; Wang et al., 2024c). The typical architecture of such an agent encompasses an LLM-based Agent, a memory module, a planning component, and a tool utilization module, as depicted in Figure 14.
@@
Figure 15. MetaGPT integrates human workflow efficiencies into LLM-based multi-agent collaboration to break down complex code-related tasks into specific, actionable procedures. These procedures are then assigned to various roles, such as Product Manager, Architect, and Engineer played by LLM. The image is sourced from the original paper (Hong et al., 2023).
In the realm of automated code generation, LLM-powered autonomous agents have demonstrated remarkable proficiency. For instance, AgentCoder (Huang et al., 2023) achieved a groundbreaking pass@1 of 96.3% on the HumanEval benchmark, forwarding a step closer to the future of automated software development (Ishibashi and Nishimura, 2024). The innovative meta-programming framework termed MetaGPT (Hong et al., 2023) integrates human workflow efficiencies into LLM-based multi-agent collaboration, as shown in Figure 15. Furthermore, (Huang et al., 2023) introduces AgentCoder, a multi-agent framework composed of three specialized agents, each with distinct roles and capabilities. These roles include a programmer agent responsible for code generation, a test designer agent tasked with generating unit test cases, and a test executor agent that executes the code and provides feedback. This division of labor within AgentCoder promotes more efficient and effective code generation. CodeAct (Wang et al., 2024a) distinguishes itself by utilizing executable Python code to consolidate LLM agent actions within a unified action space, in contrast to the generation of JSON or textual formats. Additionally, AutoCodeRover (Zhang et al., 2024b) is proposed to autonomously resolve GitHub issues for program enhancement.
@@
Despite these remarkable advancements, the journey toward fully realized AI software engineers employing LLM-powered autonomous agents is far from complete (Xi et al., 2023; Wang et al., 2024c). Critical aspects such as prompt design, context length, agent count, and toolsets call for further refinement and optimization, especially as problem complexities escalate (Ishibashi and Nishimura, 2024).
@@ 5.10. Evaluation
Table 9. The performance comparison of LLMs for code generation on the HumanEval (Chen et al., 2021) benchmark, measured by Pass@1. For models with various sizes, we report only the largest size version of each model with a magnitude of B parameters. denotes instruction-tuned models.
Model Size pass@1 (%) Availability
Closed Source GPT-4o-0513 (OpenAI, 2024a) - 91.0 [API Access]
GPT-4-Turbo-0409 (OpenAI, 2024b) - 88.2 [API Access]
GPT-4-1106 (Achiam et al., 2023) - 87.8 [API Access]
GPT-3.5-Turbo-0125 (OpenAI, 2022) - 76.2 [API Access]
Claude-3.5-Sonnet (Anthropic, 2024) - 92.0 [API Access]
Claude-3-Opus (Anthropic, 2024) - 84.9 [API Access]
Claude-3-Sonnet (Anthropic, 2024) - 73.0 [API Access]
Claude-3-Haiku (Anthropic, 2024) - 75.9 [API Access]
Gemini-1.5-Pro (Reid et al., 2024) - 84.1 [API Access]
Gemini-1.5-Flash (Reid et al., 2024) - 74.3 [API Access]
Gemini-1.0-Ultra (Reid et al., 2024) - 74.4 [API Access]
Gemini-1.0-Pro (Reid et al., 2024) - 67.7 [API Access]
PanGu-Coder2 (Shen et al., 2023) 15B 61.64 -
PanGu-Coder (Christopoulou et al., 2022) 2.6B 23.78 -
Codex (Chen et al., 2021) 12B 28.81 Deprecated
PaLM-Coder (Chowdhery et al., 2023) 540B 36 -
AlphaCode (Li et al., 2022a) 1.1B 17.1 -
Open Source Codestral (MistralAI, 2024) 22B 81.1 [Checkpoint Download]
DeepSeek-Coder-V2-Instruct (Zhu et al., 2024) 21B (236B) 90.2 [Checkpoint Download]
Qwen2.5-Coder-Instruct (Hui et al., 2024) 7B 88.4 [Checkpoint Download]
Qwen2.5-Coder (Hui et al., 2024) 7B 61.6 [Checkpoint Download]
StarCoder2-Instruct (Yuxiang Wei, 2024) 15.5B 72.6 [Checkpoint Download]
CodeGemma-Instruct (CodeGemma Team et al., 2024) 7B 56.1 [Checkpoint Download]
CodeGemma (CodeGemma Team et al., 2024) 7B 44.5 [Checkpoint Download]
StarCoder 2 (Lozhkov et al., 2024) 15B 46.3 [Checkpoint Download]
WaveCoder-Ultra (Yu et al., 2023) 6.7B 79.9 [Checkpoint Download]
WaveCoder-Pro (Yu et al., 2023) 6.7B 74.4 [Checkpoint Download]
WaveCoder-DS (Yu et al., 2023) 6.7B 65.8 [Checkpoint Download]
StableCode (Pinnaparaju et al., 2024) 3B 29.3 [Checkpoint Download]
CodeShell (Xie et al., 2024) 7B 34.32 [Checkpoint Download]
[…]
CodeQwen1.5 (Team, 2024) 7B 51.8 [Checkpoint Download]
DeepSeek-Coder-Instruct (Guo et al., 2024) 33B 79.3 [Checkpoint Download]
DeepSeek-Coder (Guo et al., 2024) 33B 56.1 [Checkpoint Download]
replit-code (Replit, 2023) 3B 20.12 [Checkpoint Download]
[…]
Magicoder-CL (Wei et al., 2023b) 7B 60.4 [Checkpoint Download]
WizardCoder (Luo et al., 2023) 33B 79.9 [Checkpoint Download]
CodeFuse (Liu et al., 2023b) 34B 74.4 [Checkpoint Download]
Phi-1 (Gunasekar et al., 2023) 1.3B 50.6 [Checkpoint Download]
Code Llama-Instruct (Roziere et al., 2023) 70B 67.8 [Checkpoint Download]
Code Llama (Roziere et al., 2023) 70B 53.0 [Checkpoint Download]
OctoCoder (Muennighoff et al., 2023) 15.5B 46.2 [Checkpoint Download]
CodeGeeX2 (Zheng et al., 2023c) 6B 35.9 [Checkpoint Download]
InstructCodeT5+ (Wang et al., 2023b) 16B 35.0 [Checkpoint Download]
CodeGen-NL (Nijkamp et al., 2022) 16.1B 14.24 [Checkpoint Download]
CodeGen-Multi (Nijkamp et al., 2022) 16.1B 18.32 [Checkpoint Download]
CodeGen-Mono (Nijkamp et al., 2022) 16.1B 29.28 [Checkpoint Download]
StarCoder (Li et al., 2023a) 15B 33.60 [Checkpoint Download]
CodeT5+ (Wang et al., 2021) 16B 30.9 [Checkpoint Download]
CodeGen2 (Nijkamp et al., 2023) 16B 20.46 [Checkpoint Download]
SantaCoder (Allal et al., 2023) 1.1B 14.0 [Checkpoint Download]
InCoder (Fried et al., 2022) 6.7B 15.2 [Checkpoint Download]
PolyCoder (Xu et al., 2022) 2.7B 5.59 [Checkpoint Download]
CodeParrot (Tunstall et al., 2022) 1.5B 3.99 [Checkpoint Download]
@@
Despite the impressive capabilities of LLMs, they exhibit a range of behaviors that are both beneficial and potentially risky. These behaviors can enhance performance across various downstream tasks but may also introduce reliability and trustworthiness concerns in LLM deployment (Chen et al., 2021; Xu et al., 2022; Chang et al., 2024). Consequently, it is imperative to develop precise evaluation approaches to discern the qualitative and quantitive differences between models, thereby encouraging further advancements in LLM capabilities.